• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name jpchen

    Log entry time 16:00:44 on February23,2009

    Entry number 262925

    keyword=Day shift summary

    Day shift summary:

    Shift workers: J. P. Chen (SL), R. Michael (TO), S. Golge (3rd)

    8:00 Continue production Kin. D, LHRS=0.9 GeV, at E=5.892 GeV,
    beam 1/2 plate OUT, target longitudinal (0 degree), 15 uA.

    8:21-8:30 found BB daq Deadtime 100%. ROC 10 not respond, rebooted and restarted CODA, all is fine.
    9:00 Completed Kin. D, total charge 0.71 C.

    9:00 Beam stopped for RF recovery/pass change 5 to 4.
    Requested controlled access for BB trigger investigation (Kalyan, Xin and Brad) and RHRS work for Parity people (Bob and Ramesh). Also
    Compton people: Alexandre, Dave Abbot, Diane and Frank. Jack.

    BB runs:
    run 1717, 7:21-8:20, 7M.
    1718, 8:21-8:24, 100% dead time, Junk
    1719, 8:31-9:03, 0.4M.

    LHRS runs:
    20354, 7:29-8:23, 0.6M
    20358, 8:32-9:03, 335K.

    9:50 changed LHRS from 0.9 GeV to 1.6 GeV

    11:00 calibrated reference cell system pressure and vaccum gauge.

    11:50 Target internal temperatue test for longitudinal pumping
    13:15 Yi/Jin: Target pumping direction changed from longitudinal to transverse+. Replaced faulty pre-amp for target chamber NMR.
    Checked COMET laser power. Found one COMET laser internal fiber
    end broken. Replaced with the last spare fiber.

    14:55 MCC called that they fixed problem in the inejctor and will
    need hall a in beam permit in 1/2 hour.

    15:30 MCC called to get people out.
    15:50 Hall in beam permit.
    15:55 filled reference cell with 114 psig N2.
    16:00 Hall in beam permit. MCC tuning beam.